OpenStudy (anonymous):

Help! Suppose for this problem that Pr[A|B]=3/8 Pr[A]=9/20 Pr[B']=[3/5] What is Pr[B|A]?

OpenStudy (anonymous):

Pr[B]=2/5 so I thought it would be (2/5)/(9/20), but it says it's wrong!

OpenStudy (anonymous):

what you are missing is only the probability of \(A\cap B\)

OpenStudy (anonymous):

\[P(B|A)=\frac{P(A\cap B)}{P(A)}\]

OpenStudy (anonymous):

I thought A∩B = Pr[B] in this problem?

OpenStudy (anonymous):

that is not usually the case in fact it would only be true if B was contained in A so that \(A\cap B=B\)

OpenStudy (anonymous):

A∩B=B∩A right? so 15/16??

OpenStudy (anonymous):

btw do not confuse sets (for example \(A\cap B\) ) with numbers like \(P(A\cap B)\) they are different animals

OpenStudy (anonymous):

\[\frac{P(A\cap B)}{P(B)}=P(A|B)\] means \[P(A\cap B)=P(A|B)P(B)\] you know both the numbers on the right, so you can compute \(P(A\cap B)\) by multiplication

OpenStudy (anonymous):

Thank you! I get it!! Thank you so much!

OpenStudy (anonymous):

you should expect that in general \(P(A\cap B)\) would be a rather small number, since it must be less than or equal to both \(P(A)\) and \(P(B)\) unlikely it is \(\frac{15}{16}\) in fact it is \[P(A\cap B)=\frac{3}{8}\times \frac{2}{5}\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Wait, so how would I find P[B|A']? Since P(A'∩B)=P(A'|B)P(B), but I don't have P[A'|B]?

OpenStudy (anonymous):

@satellite73 help please!!

OpenStudy (anonymous):

got it! nvm

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

good you know a venn diagram really helps in these if you know how to fill it in and how to visualize it

OpenStudy (anonymous):

yeah that's what I did!